# A mathematical theory for understanding when abstract representations emerge in neural networks

### v1

@@ Abstract
Recent experiments reveal that task-relevant variables are often encoded in approximately orthogonal subspaces of the neural activity space. These disentangled low-dimensional representations are observed in multiple brain areas and across different species, and are typically the result of a process of abstraction that supports simple forms of out-of-distribution generalization. The mechanisms by which such geometries emerge remain poorly understood, and the mechanisms that have been investigated are typically unsupervised (e.g., based on variational auto-encoders). Here, we show mathematically that abstract representations of latent variables are guaranteed to appear in the last hidden layer of feedforward nonlinear networks when they are trained on tasks that depend directly on these latent variables. These abstract representations reflect the structure of the desired outputs or the semantics of the input stimuli. To investigate the neural representations that emerge in these networks, we develop an analytical framework that maps the optimization over the network weights into a mean‐field problem over the distribution of neural preactivations. Applying this framework to a finite-width ReLU network, we find that its hidden layer exhibits an abstract representation at all global minima of the task objective. We further extend these analyses to two broad families of activation functions and deep feedforward architectures, demonstrating that abstract representations naturally arise in all these scenarios. Together, these results provide an explanation for the widely observed abstract representations in both the brain and artificial neural networks, as well as a mathematically tractable toolkit for understanding the emergence of different kinds of representations in task-optimized, feature-learning network models.
@@ I Introduction
How task structure shapes the geometry of neural representations has long been a central question in neuroscience and machine learning. The ability to learn appropriate representations from training data is fundamental to a system’s capacity for generalization. In neuroscience, recent experiments have shown that, following task training, neural responses across various brain regions often exhibit a characteristic low-dimensional geometry, referred to as an abstract representation [8, 17, 65, 13, 91, 24, 70, 61] (Fig. 1). In these representations, task-relevant variables are represented in distinct, approximately orthogonal subspaces in the firing rate space of neurons (Fig. 1A, right). Each of these variables is represented in an abstract format because when the activity is projected along its coding direction, it becomes invariant with respect to all the other variables (’dissociated from specific instances’). These representations are based on specialized or linearly mixed selectivity neurons and allow for a simple form of out-of-distribution generalization. Importantly, high-dimensional representations (Fig. 1B, right) do not have this property, although they would allow a linear decoder to report the value of task-relevant variables accurately [29, 44]. For these representations, the variables are nonlinearly mixed with each other [78, 29].
In machine learning, a similar type of representation geometry is called a disentangled representation [6, 93, 95, 16]. The variables represented in an abstract format in neuroscience experiments can be viewed as latent factors underlying the training data, which would be disentangled in the learned representation. Variational autoencoders and their variants have been the standard approach to obtain disentangled representations [14]. However, due to identifiability issues, it has been shown that learning disentangled representations in an entirely unsupervised manner is difficult if not impossible [35, 53, 54]. For this reason, other approaches have been proposed, which often include additional regularization [32, 3, 25, 51] or supervision [80, 40, 2].
@@
Here we focus on the supervised approach proposed in [40, 2], in which a feedforward network is trained to output multiple labels that depend on latent variables. Abstract representations of the latent variables emerge in the last hidden layer of the network after training. We study analytically the emergence of abstract representations in the simplest nonlinear network model that exhibits representation learning: a feedforward network with a single hidden layer and a nonlinear activation function (Fig. 2A, left). For the first time, we prove that the process of minimization of a simple mean square error with $l^{2}$-weight regularization in the multi-task setting of [40, 2] is guaranteed to generate abstract representations. We show that this result is robust to the choice of nonlinear activation function. To achieve his goal, we develop an analytical framework to characterize the optimal neural representation in such networks trained on any task. This analytical framework is applicable to a broader class of tasks and network architectures, providing a powerful tool for characterizing the structure of neural representation in task-optimized networks [77, 100, 81, 43, 20]. Our work therefore advances existing analytical methods for studying task-optimized neural networks, and more broadly, in those nonlinear models exhibiting permutation symmetry.
@@
The paper is organized as follows. In Section II, we define the task and network model and then introduce the analytical framework for characterizing the optimal neural representations in a two-layer nonlinear network trained on a given task. This framework establishes an exact mapping from the original network model to an effective model whose degrees of freedom are the neural preactivation patterns on training data (Fig. 2A). Solving for the optimal neural representation in the original model is then reduced to analyzing a corresponding mean-field theory in this effective model. In Section III, we apply this framework to finite-width ReLU networks and derive the optimal neural representation for the corresponding task model (Fig. 2B). In Section IV, we further use this analytical framework to show that the abstract representation remains optimal for two broad classes of nonlinear activation functions. Finally, in Section V, we extend the framework to study additional tasks and deep network architectures.
@@ II.1 Data and network model
Figure 1: Abstract representations. (A) Each stimulus input in the task (e.g. images of handwritten digits) is associated with several binary variables (e.g. parity and magnitude of the digits). An abstract representation is one in which each binary variable is represented along a single axis in the population activity space. This is shown in the top plot inside the frame. This geometric property can be quantified using the parallelism score (PSPS), which measures how parallel the coding directions for one variable remain when the other variables vary. For example, it would measure the parallelism of the parity coding direction for small and large digits (two values of the other variable, magnitude). Abstract representations are low-dimensional and have $PS=1$. In an alternative neural representation, the points representing the different digits are arranged on a tetrahedron shape, which is the highest dimensional representation (bottom in the frame). This would correspond to a non-abstract representation and has $PS\sim 0$. (B) Magnitude and parity are equally decodable for both geometries.
Here we assume that the input patterns are basically unstructured, whereas the interesting structure is in the desired outputs, or the labels used to train the network (similarly to [83, 40]). In general, these labels might reflect the structure of a latent space used to generate the complex, seemingly unstructured inputs [40]. However, here we ignore this possibility and we just assume that all the structure that we are going to study is in the output patterns, $\mathbf{y}^{i}$.
More specifically, we assume that the output associated with a particular input $\mathbf{y}^{i}\in\{\pm 1\}^{d_{Y}}$ consists of $d_{Y}\in\mathbb{N}_{+}$ binary labels. We denote the input and output matrices as
@@
Based on their output labels, all the training data form $2^{d_{Y}}$ distinct classes. We further assume that each class has the same number of training data, $n\in\mathbb{N}_{+}$. Namely, all the classes are balanced, and the total number of training data points $P$ satisfies
@@
Denote the $i$th row vector of $Y$ as $\mathbf{v}_{i}\in\{\pm 1\}^{P}$, whose components are the $i$th binary labels for all training data, $i=1,2,..d_{Y}$. Under the above assumption, we find (SI §1.1) that (i) $\mathbf{v}_{i}\cdot\mathbf{v}_{j}=P\delta_{ij}$; (ii) $\mathbf{1}\cdot\mathbf{v}_{i}=0$, for $\mathbf{1}=(1,1,...,1)^{T}\in\mathbb{\mathbb{R}}^{P}$, i.e. exactly half of the components of $\mathbf{v}_{i}$ is +1+1 and the other half is 1-1, and (iii) $\mathbf{v}_{i}$’s are all the eigenvectors of the output kernel matrix $K_{Y}\equiv Y^{T}Y$ with non-vanishing eigenvalues.
@@
$=$ $\sum_{i=1}^{P}\left[\mathbf{y}^{i}-W_{2}\phi(W_{1}\mathbf{x}^{i}+\mathbf{b})\right]^{2}+\lambda_{1}\left\|W_{1}\right\|_{F}^{2}$
$\hskip 90.00014pt+\lambda_{1}\left\|\mathbf{b}\right\|_{2}^{2}+\lambda_{2}\left\|W_{2}\right\|_{F}^{2}$
@@ II.2 Parallelism score measures the abstractness of the neural representation
For different labels $\alpha_{1},\alpha_{2}\in\{\pm 1\}^{d_{Y}-1}$, we quantify how consistent the direction of representation changes are, using cosine similarity
@@
The parallelism score for the $k$th latent label is defined as the average cosine similarity for all distinct pairs $\left(\alpha_{1},\alpha_{2}\right),$
@@
By definition, $PS$ of a neural representation only depends on the inner product between the neural representation of all data pairs, $\mathbf{r}^{i}\cdot\mathbf{r}^{j}$. So we introduce the representation kernel matrix $K\in\mathbb{R}^{P\times P}$ whose element is
@@
This kernel matrix fully determines the $PS$ for a neural representation. Moreover, following the definition, $PS$ does not change when (1) adding a constant to all elements of $K$, or (2) multiplying $K$ by a positive number.
If the neural representation of data is completely random, $PS\sim 0$. A neural representation is called an abstract representation if the PSPS is large and close to 11 [8, 17]. We emphasize that the $PS$ of a neural representation is not intrinsically tied to the (linear) decodability of the latent labels (Fig. 1A).
@@ II.3 The effective mean-field energy
Since up to a rotation, the input and output data $X,Y$ can be reconstructed from these kernel matrices, we say that these kernel matrices capture the input and output geometry (Fig. 2B). Denote $\text{Range}K_{X}$ as the column space of the matrix $K_{X}$
@@
where the $k$th column vector $\mathbf{\mathbf{h}}_{k}\in\mathbb{R}^{P}$ represents the preactivation pattern of the $k$th hidden neuron for all $P$ data points (or stimulus conditions). We note that $H$ is closely related to the response matrix commonly used in neuroscience [44], whose rows correspond to all the stimulus conditions and columns to all the neurons.
Figure 2: Model set-up and summary of the main results. (Middle) The two-layer nonlinear network models are trained on tasks related to abstract representation. The weight and bias parameters in each layer are optimized for the task. The hidden layer has width $M$. For a range of input geometries (characterized by different input kernel matrices $K_{X}$) and the specified output geometry (where each output label is exactly the latent label), the optimal hidden representation is always abstract. (Top right) The output labels for each stimulus input are exactly its associated binary latent labels. (Top left) The range of input geometry changes smoothly from a fully orthogonalized input where different stimuli are represented by orthogonal vectors, to an input geometry that is fully aligned with the outputs. Here to illustrate the orthogonalized input in the 44-dimensional space, we draw the 3D projection of it that has a tetrahedron shape.
With these notations, we find that (SI §1.2) the zero-temperature free energy can be obtained via the following optimization problem over the neural preactivations,
@@
Here $K_{X}^{\dagger}$ is the Moore-Penrose inverse of $K_{X}$. Eq. (12)-(II.3) shows that the zero-temperature free energy is determined by the global minima of the effective energy function [Eq. (II.3)] over the neural preactivation patterns in the hidden layer, subject to the constraint $\mathbf{h}_{k}\in\text{Range}K_{X}$.
This maps the original system, whose energy function [Eq. (II.1)] is over its parameter space $(W_{1},W_{2},\mathbf{b})$, into an effective system with energy function described by Eq. (II.3). The effective system consists of $M$ neurons, each of which has a $P$-dimensional state variable $\mathbf{h}\in\mathbb{R}^{P}$, lying in the data-dependent subspace $\text{Range}K_{X}$ (Fig. 3DE). Note that Eq. (12)-(II.3) are valid for any network width $M$ and input/output kernel matrix $K_{X}/K_{Y}$.
@@
Eq.(II.3) is invariant under permutation over neurons. So we introduce the empirical measure of the preactivations $\rho_{M}=\sum_{k=1}^{M}\delta_{\mathbf{h}_{k}}$ (the unnormalized empirical distribution of $\mathbf{h}_{k}$’s) and rewrite it as
@@
Eq. (14) shows that the energy function of the effective $M$-neuron system only depends on the global statistics of neural activity $\rho_{M}$. The measure $\rho_{M}$ can be regarded as the order parameter of the (finite-size) effective $M$-neuron system, which is permutation-invariant and reflects the symmetry in the original system [Eq. (II.1)]. This is similar to many other models in statistical physics [1, 47, 46, 63, 42, 69] where the order parameter is a (probability) measure rather than simple scalars.

### v2

@@ Abstract
Recent experiments in neuroscience reveal that task-relevant variables are often encoded in approximately orthogonal subspaces of neural population activity. These disentangled, or abstract, representations have been observed in multiple brain areas and across different species. These representations have been shown to support out-of-distribution generalization and rapid learning of novel tasks. The mechanisms by which these representations emerge remain poorly understood, especially in the case of supervised task behavior. Here, we show mathematically that abstract representations of latent variables are guaranteed to appear in the hidden layer of feedforward nonlinear networks when they are trained on tasks that depend directly on these latent variables. These learned abstract representations reflect the semantics of the input stimuli. To show this, we reformulate the usual optimization over the network weights into a mean‐field optimization problem over the distribution of neural preactivations. We then apply this framework to finite-width ReLU networks and show that the hidden layer of these networks will exhibit an abstract representation at all global minima of the task objective. Finally, we extend our findings to two broad families of activation functions as well as deep feedforward architectures. Together, our results provide an explanation for the widely observed abstract representations in both the brain and artificial neural networks. In addition, the general framework that we develop here provides a mathematically tractable toolkit for understanding the emergence of different kinds of representations in task-optimized, feature-learning network models.
@@ I Introduction
How task structure shapes the geometry of neural representations has long been a central question in neuroscience and machine learning. The ability to learn appropriate representations from training data is fundamental to a system’s capacity for generalization. In neuroscience, recent experiments have shown that, following task training, neural responses across various brain regions often exhibit a characteristic low-dimensional geometry, referred to as an abstract (or, in machine learning, a disentangled) representation [7, 15, 64, 12, 90, 23, 65, 61] (Fig. 1). In these representations, task-relevant variables are represented in distinct, approximately orthogonal subspaces in the firing rate space of neurons (Fig. 1A, right). One possibility is that these subspaces correspond to distinct sets of neurons. In this case, a single neuron would represent the value of a single variable. The discovery of an unsupervised learning algorithm that will produce this kind of ”disentangled” representation has been a topic of great interest in machine learning [30, 58, 24]. Another possibility is that these subspaces are not necessarily aligned with single neurons, but instead distributed across the activity of the entire population of neurons, as seen in many experimental work from neuroscience [7, 15, 23], where all neurons in the population play at least a small part in representing almost all variables. Whether expressed in specialized sub-populations or across the whole population, abstract representations make the representation of a single variable invariant to all other variables (’dissociated from specific instances’). As a consequence, they facilitate certain forms of out-of-distribution generalization and rapid learning of novel tasks. In contrast, non-abstract representations (Fig. 1B, right) – in which different variables are not invariant to each other – do not have these properties, though the relevant variables can still be accurately decoded [27, 43]. In these non-abstract representations, the different variables are nonlinearly mixed with each other [74, 27] – producing a high-dimensional representation that can be used to decode any stimulus condition dichotomy.
In machine learning, variational autoencoders and their variants have been the standard approach to obtain disentangled representations [13]. However, due to identifiability issues, it has been shown that learning disentangled representations in an entirely unsupervised manner is difficult if not impossible [34, 52, 53]. For this reason, other approaches have been proposed, which often include additional regularization [31, 2, 24, 50] or supervision [76, 39, 1]. These supervised approaches have a clear connection to neuroscience: through natural behavior, humans and other animals learn to interact with and classify stimuli in their environment in a variety of ways. Previous work has shown that learning to classify single stimuli in multiple ways can naturally give rise to abstract representations in artificial neural networks [39]. Thus, abstract representations may emerge in the brain through this behavioral process. However, this prior work did not establish general conditions in which these representations emerge, nor did they develop a mathematical theory that explains this emergence.
@@
Here, we formulate an analytical theory that explains the emergence of abstract representations in feedforward artificial neural networks trained to perform multiple tasks that share the same set of latent variables [39, 1]. We begin by studying the emergence of abstract representations in the simplest nonlinear network model that exhibits representation learning: a feedforward network with a single hidden layer and nonlinear activation functions (Fig. 2A, left). For the first time, we prove that the process of minimization of a simple mean square error with $l^{2}$-weight regularization in the multi-task setting of [39, 1] is guaranteed to generate abstract representations. We show that this result is robust to the choice of nonlinear activation function. However, our analytical framework can be applied well beyond the tasks studied here. In particular, it can be used to characterize the optimal neural representation in nonlinear feature-learning neural networks trained on any task. Our framework provides a powerful tool for characterizing the structure of neural representation in task-optimized networks [73, 99, 82, 42, 18]. Our work therefore advances existing analytical methods for studying task-optimized neural networks, and more broadly, in those nonlinear models exhibiting permutation symmetry.
@@
The paper is organized as follows. In Section II, we define the task and network model and then introduce the analytical framework for characterizing the optimal neural representations in a two-layer nonlinear network trained on a given task. This framework establishes an exact mapping from the original network model to an effective model whose degrees of freedom are the neural preactivation patterns on training data (Fig. 3A). Solving for the optimal neural representation is then reduced to a corresponding mean-field theory of this effective model. In Section III, we apply this framework to finite-width ReLU networks and derive the optimal neural representation for the corresponding task model (Fig. 3B). In Section IV, we further use this analytical framework to show that the abstract representation remains optimal for two broad classes of nonlinear activation functions. Finally, in Section V, we extend the framework to study additional tasks and deep network architectures.
@@ II.1 Data and network model
Figure 1: Abstract representations. (A) Each stimulus input in the task (e.g. images of handwritten digits) is associated with several binary latent variables (e.g. parity and magnitude of the digits). These latent variables are exactly the output labels $\mathbf{y}^{i}$ in our data model Eq. (1). An abstract representation is one in which each binary variable is represented along a single axis in the population activity space, shown in the top plot inside the frame. This geometric property can be quantified using the parallelism score (PSPS), which measures how parallel the coding directions for one variable remain when the other variables vary. For example, it would measure the parallelism of the parity coding direction for small and large digits (corresponding to two values of the other variable, magnitude). Abstract representations are low-dimensional and have $PS=1$. In an alternative neural representation, the points representing the different digits are arranged on a tetrahedron shape, which is the highest-dimensional representation (bottom in the frame). This would correspond to a non-abstract representation and has $PS\sim 0$. (B) Magnitude and parity are equally linearly decodable for both geometries.
We assume that the input patterns are unstructured (that is, uncorrelated with each other), whereas the desired outputs (i.e., the training labels) have low-dimensional structure (similar to [81, 39]). In general, these labels might reflect the structure of a latent space used to generate the complex, seemingly unstructured inputs and might depend on the latent variables in a complex way [39]. Here, we assume that the network’s desired outputs, $\mathbf{y}^{i}$, are exactly these latent variables, and ignore the possibility of a more complex dependency between the latent variables and output labels. An example of such dataset is the MNIST images where the latent variables are magnitude and parity (Fig. 1A).
More specifically, the output associated with a particular input $\mathbf{y}^{i}\in\{\pm 1\}^{d_{Y}}$ is assumed to consist of $d_{Y}\in\mathbb{N}_{+}$ binary labels. In the typical multitask set-up [76], different binary labels represent different ways to classify the inputs, usually based on different features of the inputs (Fig. 1). We define the input and output data matrices as
@@
Based on their output labels, all the training data form $2^{d_{Y}}$ distinct classes. We further assume that each class has the same number of training data, $n\in\mathbb{N}_{+}$. Namely, all the classes are balanced, and thus the total number of training data points $P$ satisfies
@@
Denote the $i$th row vector of $Y$ as $\mathbf{v}_{i}\in\{\pm 1\}^{P}$, whose components are the $i$th binary labels for all training data, $i=1,2,..d_{Y}$. Under the above assumption, we find that (i) $\mathbf{v}_{i}\cdot\mathbf{v}_{j}=P\delta_{ij}$, where $\delta_{ij}$ is the Kronecker delta notation; (ii) $\mathbf{1}\cdot\mathbf{v}_{i}=0$, for $\mathbf{1}=(1,1,...,1)^{T}\in\mathbb{\mathbb{R}}^{P}$, i.e. exactly half of the components of $\mathbf{v}_{i}$ is +1+1 and the other half is 1-1, and (iii) $\mathbf{v}_{i}$’s are all the eigenvectors of the output kernel matrix $K_{Y}\equiv Y^{T}Y$ with non-vanishing eigenvalues (SI §1.1).
@@
$=$ $\sum_{i=1}^{P}\left[\mathbf{y}^{i}-W_{2}\phi(W_{1}\mathbf{x}^{i}+\mathbf{b})\right]^{2}$
$\hskip 37.00002pt+\lambda_{1}\left\|W_{1}\right\|_{F}^{2}+\lambda_{1}\left\|\mathbf{b}\right\|_{2}^{2}+\lambda_{2}\left\|W_{2}\right\|_{F}^{2}$
@@ II.2 Parallelism score measures the abstractness of the neural representation
For different labels $\alpha_{1},\alpha_{2}\in\{\pm 1\}^{d_{Y}-1}$, we quantify how consistent the direction of representation changes are when varying the $k$th label, using cosine similarity
@@
The average parallelism score for the $k$th latent label is defined as the average cosine similarity for all distinct pairs $\left(\alpha_{1},\alpha_{2}\right),$
@@
By definition, $PS$ of a neural representation only depends on the inner product between the neural representation of all data pairs, $\mathbf{r}^{i}\cdot\mathbf{r}^{j}$. So we define the representation kernel matrix $K\in\mathbb{R}^{P\times P}$ whose element is
@@
This kernel matrix fully determines the $PS$ for a neural representation. Moreover, following the definition, $PS$ does not change when (1) adding a constant to all elements of $K$, or (2) multiplying $K$ by a positive scalar.
If the neural representation of data is completely random, $PS\sim 0$. A neural representation is called an abstract representation if the $PS$ is large and close to $1$ [7, 15]. However, even a representation with $PS\sim 0$ allows for a linear decoder to recover the latent labels (Fig. 1A). Thus, the emergence of representations with high $PS$ is not a direct consequence of high task performance.
@@ II.3 The effective mean-field energy
Since up to a rotation, the input and output data $X,Y$ can be reconstructed from these kernel matrices, we say that these kernel matrices capture the input and output geometry (Fig. 2). Denote $\text{Range}K_{X}$ as the column space of the matrix $K_{X}$
@@
where the $k$th column vector $\mathbf{\mathbf{h}}_{k}\in\mathbb{R}^{P}$ represents the preactivation pattern of the $k$th hidden neuron for all $P$ data points (or stimulus conditions). We note that $H$ is related to the response matrix commonly used in neuroscience [43], whose rows correspond to all the stimulus conditions and columns to all the neurons.
Figure 2: Model set-up and summary of the main results. (Middle) The two-layer nonlinear network models are trained on tasks related to abstract representation. The weight and bias parameters are optimized for the task. The hidden layer has width $M$. For a range of input geometries (characterized by different input kernel matrices $K_{X}$) and the specified output geometry (where each output label is exactly the latent label), the optimal hidden representation is abstract. (Top right) The output labels for each stimulus input are exactly its associated binary latent labels. (Top left) The range of input geometry smoothly transitions from a fully orthogonalized input, where different stimuli are represented by orthogonal vectors, to a fully output-aligned input (where $K_{X}\approx K_{Y}$). To illustrate the fully orthogonalized input in the 44-dimensional space, we draw its 3D projection, which forms a tetrahedron.
We find that (SI §1.2) the zero-temperature free energy can be obtained via the following optimization problem over the neural preactivations,
@@
Here $K_{X}^{\dagger}$ is the Moore-Penrose inverse of $K_{X}$. Eqs. (12)-(II.3) shows that the zero-temperature free energy is determined by the global minima of the effective energy function [Eq. (II.3)] over the neural preactivation patterns in the hidden layer, subject to the constraint $\mathbf{h}_{k}\in\text{Range}K_{X}$.
This result maps the original system, whose energy function [Eq. (II.1)] is over its parameter space $(W_{1},W_{2},\mathbf{b})$, into an effective system with energy function described by Eq. (II.3). The effective system consists of $M$ neurons, each of which has a $P$-dimensional state variable $\mathbf{h}\in\mathbb{R}^{P}$, lying in the data-dependent subspace $\text{Range}K_{X}$ (Fig. 3DE). Note that Eqs. (12)-(II.3) are valid for any network width $M$ and input/output kernel matrix $K_{X}/K_{Y}$.
@@
Eq.(II.3) is invariant under permutation over neurons (note that the $h_{k}$ only appear in summation). So we introduce the empirical measure of the preactivations $\rho_{M}=\sum_{k=1}^{M}\delta_{\mathbf{h}_{k}}$ (the unnormalized empirical distribution of $\mathbf{h}_{k}$’s) and rewrite the sums as integrals over this new distribution, so
@@
Eq. (14) shows that the energy function of the effective $M$-neuron system only depends on the global statistics of neural activity $\rho_{M}$. Thus, the measure $\rho_{M}$ can be viewed as the order parameter of the (finite-size) effective $M$-neuron system, which is permutation-invariant and reflects the symmetry in the original system [Eq. (II.1)]. This is similar to many other models in statistical physics [91, 45, 46, 62, 41, 66] where the order parameter here is a (probability) measure rather than simple scalars.